3x+4y=12
(0, ), ( ,3/4),( ,0), (8/3, )
Complete the ordered pair so that each is a solution for the given equation.

Answer by stanbon(75887)

3x+4y=12
(0, ), ( ,3/4),( ,0), (8/3, )
Complete the ordered pair so that each is a solution for the given equation.
---------------------
If x=0,4y=12 and y=3
If y=(3/4), 3x+3=12, x=3
If y=0, 3x=12, x=4
If x=(8/3), 3*(8/3)+4y=12,4y=4, y=1
===============
Cheers,
Stan H.